the average weight of 5-man Freshman basketball team was thought to be 60 kg. However, a mistake was made-- one of the weights was written down wrong. Once the mistake was found, and that one weight was changed to the right value of 80 kg, the correct average turned out to be 70 kg. What was the mistaken value? That is, what was written down at first instead of 80 kg?
Answer by mananth(16949)

the average weight of 5-man Freshman basketball team was thought to be 60 kg. However, a mistake was made-- one of the weights was written down wrong. Once the mistake was found, and that one weight was changed to the right value of 80 kg, the correct average turned out to be 70 kg. What was the mistaken value? That is, what was written down at first instead of 80
...
let the weight noted be x kg?
..
5 men .... average = 60...... total weight = 300 kg.
...
(300 -x +80)/5 = 70
380-x= 350
add -380
-x=350-380
-x=-30
/-1
x= 30 this was the weight recorded earlier.
...
m.ananth@hotmail.ca